if one side of a square is increased by 8 feet and an adjacent side is decreased by 2 feet a rectangle is formed whose perimeter is 40 feet. find the length of a side of the square
Found 2 solutions by jorel1380, htmentor:
Answer by jorel1380(3719)

2(n+8)+2(n-2)=40
2n+16+2n-4=40
4n=28
n=7
The square is 7 feet by 7 feet..

Answer by htmentor(1343)

Let s = the length of a side of the square
The two new sides have length s + 8 and s - 2
The equation for the new perimeter is:
2(s+8) + 2(s-2) = 40
Solve for s:
2s + 16 + 2s - 4 = 40
4s = 28
s = 7